Calculate Ea for a reaction having k = 2.61 x 10-5 at 190.0°C, and k = 3.02 x 10-3 at 250.0°C.

a. 2.56 x 105

b. 1.59 x 105

c. 39.5

d. 3.13 x 104
Answer #1

in the ( ) = 10*** Ti190+273 = 463k T2 =250+213 = 523K moliy - 2 in l. 157088 X10 Ea 160 = - 8. 314x103 Kg/mol (242 EO 4.75 =